# Mind the Gap: The Divergence Between Human and LLM-Generated Tasks

### v2

@@ Abstract
Humans constantly generate a diverse range of tasks guided by internal motivations. While generative agents powered by large language models (LLMs) aim to simulate this complex behavior, it remains uncertain whether they operate on similar cognitive principles. To address this, we conducted a task-generation experiment comparing human responses with those of an LLM agent (GPT-4o). We find that human task generation is consistently influenced by psychological drivers, including personal values (e.g., Openness to Change) and cognitive style. Even when these psychological drivers are explicitly provided to the LLM, it fails to reflect the corresponding behavioral patterns. They produce tasks that are markedly less social, less physical, and thematically biased toward abstraction. Interestingly, while the LLM’s tasks were perceived as more fun and novel, this highlights a disconnect between its linguistic proficiency and its capacity to generate human-like, embodied goals. We conclude that there is a core gap between the value-driven, embodied nature of human cognition and the statistical patterns of LLMs, highlighting the necessity of incorporating intrinsic motivation and physical grounding into the design of more human-aligned agents.
@@ Introduction
Humans can generate an infinite variety of goals to guide their behavior and enrich their daily lives, often infused with a distinct personal touch. This ability of autonomous task generation plays a central role in human cognition, shaping how individuals adapt to and interact with the world (Chu, Tenenbaum, and Schulz 2024; Molinaro and Collins 2023). Recent advances in large language models (LLMs) have enabled generative agents that simulate human activity in virtual worlds with striking autonomy (Park et al. 2023; Yang et al. 2024). These agents can initiate tasks, interact with others, and maintain activity over time, often producing goal-directed, human-like behavior. Yet the core question remains: does this simulation capture the deep cognitive mechanisms that drive human behavior, or is it merely a sophisticated mimicry of its surface patterns? Can LLMs, trained on vast corpora of text, truly replicate the diversity, personal flair, and intrinsic motivation that characterize human goal-setting?
@@
Second, human goals are shaped by embodied experience, that is, by the sensorimotor constraints of the body and accumulated interactions with the physical and social world (Varela, Thompson, and Rosch 2017; Foglia and Wilson 2013; Xiang et al. 2023). Humans naturally understand objects not just by semantic labels, but by how they can be physically manipulated and used in context. This grounding shapes how we imagine, evaluate, and select possible actions in real-world settings.
@@
To systematically investigate these issues, we designed a novel text-based task generation paradigm that elicits unconstrained, intrinsically motivated responses from both humans and LLMs (Figure 1). Our approach proceeds in two steps.
@@
Second, we compare this human baseline to the output of a widely used LLM (GPT-4o), conditioned on the same value. This allows us to evaluate whether such models can replicate the psychological signatures of autonomous goal-setting or whether fundamental differences remain, particularly in terms of motivational grounding and embodied realism.
@@
Recent advances have positioned LLMs as powerful tools for simulating human cognition and behavior. Studies have demonstrated that LLM-based “generative agents” can navigate complex environments, autonomously generate plans, and engage in social interactions that exhibit emergent dynamics (Park et al. 2023; Yang et al. 2025). These capabilities have led to growing interest in using LLMs as cognitive models to explore how humans think, decide, and act.
Despite these capabilities, researchers have identified significant gaps in the psychological plausibility of these models. While LLMs can mimic surface-level reasoning and planning, they often lack deeper cognitive mechanisms such as emotion, causality, physical dynamics and social cognition (Hu, Sosa, and Ullman 2025; Binz and Schulz 2023; Ullman 2024). Moreover, because they are trained purely on text, LLMs lack the sensorimotor grounding that supports embodied goal formation in humans. Embodied cognition theories emphasize that human goals and reasoning are deeply shaped by our physical experiences and social embeddedness (Varela, Thompson, and Rosch 2017). This suggests that generating goals rooted in physical interaction and embodied experience is a critical test case for the fidelity of LLM simulators—a test we conduct in this study.
@@ Methods
We conducted two experiments to identify the unique signatures of human task generation by contrasting it with the output of a large language model (LLM). Experiment 1 elicited tasks from both human participants and an LLM agent. For the human baseline, we measured core psychological constructs hypothesized to drive autonomous behavior: stable motivational drivers (personal values, via PVQ21, Schwartz 2021), cognitive strategies for navigating complexity (thinking style, via TWS, Sagiv et al. 2010), along with emotional states as controls. Experiment 2 then involved an independent sample of human raters who evaluated key attributes of the generated tasks, enabling a direct comparison between human and AI outputs.
@@
The experiment employed a 2 (environmental context: high vs. low complexity) $\times$ 2 (social context: social vs. non-social) between-subjects factorial design (Figure 2B)  [^2] Experiment link: https://value-task-collect.pages.dev/consent. The environmental context was manipulated to vary the creative constraints. In the high-complexity condition, participants were presented with a list of 32 common living room items. To increase the cognitive demand for task generation in the low-complexity condition, we removed seven items with high entertainment affordances (e.g., Dice, Basketball), leaving 25 common items (e.g., Pens, Chairs).
@@
All object stimuli were selected from the room asset library of the virtual simulation based on Unreal Engine 5 platform  [^3] https://www.unrealengine.com/en-US/unreal-engine-5 for ecological validity.
@@
After completing the psychological scales, participants were randomly assigned to one of the four conditions. They were shown a list of objects and instructed: “Please imagine that you are in a real room with several items in it. You need to use these items to pass the time.” For each task, participants provided a title, a list of used objects, a description, task setup, goal, and scoring system (Figure 2A). They reported the mental effort and perceived difficulty for each task generated. Participants were required to generate at least three tasks before choosing to conclude the experiment.
LLM Agent Implementation
We implemented two GPT-4o agent conditions to test whether the model’s deviation from human goal-setting stems from a simple information deficit or a fundamental mechanism differences. In the raw condition, the model received only the task prompt, serving as a baseline for its default behavior. In the matched condition, each agent was given the full psychological profile of a corresponding human participant, including values and cognitive style. This condition directly tested whether the model could translate these psychological constructs into value-driven behavior when provided as semantic data, or whether its lack of internal motivation and embodied grounding imposes a deeper constraint. Both agents operated at a temperature of 1.0 to encourage diverse outputs.
@@
The stimulus pool for evaluation comprised two sets of tasks. To capture the full spectrum of human creativity, we included all 575 valid tasks generated by human participants. To create a focused and stringent comparison, we compared this to the output from the matched GPT condition, including the first task generated for each of the 176 unique participant profiles. This ensures a representative, non-redundant sample of the model’s performance when conditioned on individual data. Tasks from the raw GPT agent were reserved for computational analyses (e.g., topic modeling) but excluded from human rating to streamline the evaluation process. Each of the 75 evaluators then rated a randomized subset of 50 tasks, yielding approximately five independent ratings per task.
@@
Evaluators rated each task on a 0-10 scale across several dimensions selected to capture creativity (Fun, Novelty), and embodiment (Mental Demand, Physical Demand, Body Part Engagement). These metrics provide a quantitative basis for comparing the qualitative nature of human- and AI-generated goals (Figure 2C).
@@
Task Content Analysis
To identify emergent themes in the generated tasks, we performed topic modeling using BERTopic (Grootendorst 2022) on the full corpus of 1718 tasks (575 human tasks, 575 matched GPT tasks, and 568 raw GPT tasks). To measure the creativity of each participant’s output, we also computed their task diversity score, defined as the mean cosine distance between the sentence-transformer embeddings (Reimers and Gurevych 2019) of the tasks they generated.
Statistical Modeling
Figure 3: Personal values and cognitive style shape human goal generation. (A) Regression coefficients of the key predictors on task attributes (Novelty, Fun, and Task Diversity). (B) Openness to Change values predict task fun and novelty. (C) Cognitive style (TWS) interacts with environment to predict task diversity. Bars indicate 95% confidence intervals. * indicates $p<0.05$, ** indicates $p<0.01$.
@@
To test our hypotheses about the link between psychological traits and task attributes, we specified a series of linear mixed-effects models (LMMs, Brauer and Curtin 2018) to predict key dependent variables: the rated attributes of Fun, Novelty, and their task content diversity. We first specified a baseline model including only demographic variables and random intercepts for participants. Then we specified a main-effects model by adding experimental conditions and psychological predictors like personal values and thinking style. Finally, we additionally included the pre-specified TWS × Environmental Context interaction into the model. The models were compared using a likelihood ratio test (LRT, Moreira 2003) to determine if including the psychological predictors significantly improved model fit.
@@ Results
Our analyses proceed in two stages. First, we establish a behavioral signature of human goal generation, demonstrating its systematic link to personal values and environmental context. Second, we contrast this human baseline against the output of LLM agents to reveal fundamental mechanism gaps.
@@
To establish a human baseline, we first tested how psychological traits predicted the goals people generated. Linear mixed-effects models confirmed that task generation was systematically linked to stable personal values. The value of Openness to Change significantly predicted higher ratings of both task novelty ($b=0.152$, $t=2.406$, $p=0.017$) and fun ($b=0.146$, $t=2.316$, $p=0.022$). In contrast, the self-interest value dimension showed no such relationship. This finding supports the core hypothesis that stable motivational principles guide autonomous goal-setting.
Behavior was not only driven by stable values but was also adapted to the environment. We found that the creativity of generated tasks depended on a significant interaction between an individual’s cognitive style (TWS) and the complexity of the environment. This pattern appeared in both subjective ratings of task fun (interaction of TWS and Environmental Context: b=0.140b=-0.140, t=2.640t=-2.640, $p=0.009$), and in objective measures of task diversity (b=0.215b=-0.215, t=2.809t=-2.809, $p=0.006$). Specifically, individuals with a more systematic thinking style generated more fun tasks in the simple environment, whereas those with an intuitive style thrived in the complex one. This suggests people flexibly deploy different cognitive strategies to match environmental demands.
Figure 4: Human and GPT generated systematically different goals. (A) Thematic distribution of human and LLM tasks. 1718 tasks generated by human and GPT were clustered into 12 topics and 3 themes: “Physical & Sports Activities”, “Relaxation & Household Activities”, and “Mental & Artistic Activities”. (B) LLMs exhibit a lower propensity for social task generation. (C) LLM tasks are perceived as more mentally and less physically demanding. Black dots stands for the overall difficulty of tasks. Orange and green dots stands for mental and physical load respectively. (D) Engagements of different body parts.
@@
Having established a human behavioral signature, we next compared it to the output of LLM agents to identify structural differences. The comparison revealed systematic differences in task content and structure.
@@
We directly compared the distribution of task topics between human participants and LLM agents. Topic modeling identified three dominant themes across the dataset: Physical & Sports, Relaxation & Household, and Mental & Artistic. The distribution of tasks across these themes differed significantly between humans and the LLM agents ($\chi^{2}(4)=581.45$, $p<0.001$). Human participants generated a balanced mix of activities. In contrast, LLMs showed a bias toward abstract tasks like music and writing (Figure 4A).
@@
Beyond thematic content, the models generated significantly fewer social tasks than humans in settings where another person was present ($\chi^{2}(2)=93.19$, $p<0.001$). In these scenarios, 58% of human-generated tasks were multiplayer, with 83% of participants creating at least one social goal. Humans proposed a range of interactive tasks, from simple conversation games like “Guess Who I Am” to competitive games like “Pillow Fight”. In contrast, multiplayer tasks from the Raw GPT and Matched GPT were far less common, accounting for only 29% and 24% of their outputs, respectively (Figure 4B).
@@ A Disembodiment Gap in Perceived Demands
The thematic differences between human and LLM tasks were reflected in their perceived mental and physical demands (Figure 4C). Tasks generated by the Matched GPT were rated as significantly more mentally demanding (z=11.64z=-11.64, FDR-corrected $p<0.001$) and less physically demanding (z=8.40z=-8.40, FDR-corrected $p<0.001$) than those generated by humans. This aligns with the LLM’s focus on abstract, “Mental & Artistic” activities.
This mental-physical divide was also evident in the specific body parts required for task execution (Figure 4D). GPT tasks demanded greater brain engagement (z=11.40z=-11.40, FDR-corrected $p<0.001$) but significantly less involvement of the arms (z=5.43z=-5.43, FDR-corrected $p<0.001$) and lower body (z=7.93z=-7.93, FDR-corrected $p<0.001$). Engagement of the eyes, ears, and hands did not differ between the two groups.
@@
Despite their disconnection from physical and social experience, Matched GPT tasks were rated as more novel (z=5.63z=-5.63, FDR-corrected $p<0.001$) and more fun (z=2.73z=-2.73, FDR-corrected $p=0.010$) than those created by humans (Figure 5). This highlights a key tradeoff: LLMs can generate text that seems creative and engaging, but they do so without grounding in embodied or socially situated experience.

### v3

@@ Abstract
Humans constantly generate a diverse range of tasks guided by internal motivations. While generative agents powered by large language models (LLMs) aim to simulate this complex behavior, it remains uncertain whether they operate on similar cognitive principles. To address this, we conducted a task-generation experiment comparing human responses with those of an LLM (GPT-4o). We find that human task generation is consistently influenced by psychological drivers, including personal values (e.g., Openness to Change) and cognitive style. Even when these psychological drivers are explicitly provided to the LLM, it fails to reflect the corresponding behavioral patterns. They produce tasks that are markedly less social, less physical, and thematically biased toward abstraction. Interestingly, while the LLM’s tasks were perceived as more fun and novel, this highlights a disconnect between its linguistic proficiency and its capacity to generate human-like, embodied goals. We conclude that there is a core gap between the value-driven, embodied nature of human cognition and the statistical patterns of LLMs, highlighting the necessity of incorporating intrinsic motivation and physical grounding into the design of more human-aligned agents.
Codehttps://github.com/Yilong-Lu/Mind˙the˙Gap
@@ Introduction
Humans can generate an infinite variety of goals to guide their behavior and enrich their daily lives, often infused with a distinct personal touch. This ability of autonomous task generation plays a central role in human cognition, shaping how individuals adapt to and interact with the world (Chu et al. 2024; Molinaro and Collins 2023). Recent advances in Large Language Models have enabled generative agents that simulate human activity in virtual worlds with striking autonomy (Park et al. 2023; Yang et al. 2024). This has spurred the development of benchmarks like TaskBench (Shen et al. 2024) and AgentGen (Hu et al. 2024), which evaluate the planning and task automation capabilities of these models. However, these performance-focused evaluations often overlook a more fundamental question: does this simulated autonomy capture the deep cognitive mechanisms that drive human behavior, or is it merely a sophisticated mimicry of surface patterns? Can LLMs, trained on vast corpora of text, truly replicate the diversity, personal flair, and intrinsic motivation that characterize human goal-setting?
@@
Second, human goals are shaped by embodied experience, that is, by the sensorimotor constraints of the body and accumulated interactions with the physical and social world (Varela et al. 2017; Foglia and Wilson 2013; Xiang et al. 2023). Humans naturally understand objects not just by semantic labels, but by how they can be physically manipulated and used in context. This grounding shapes how we imagine, evaluate, and select possible actions in real-world settings.
@@
To investigate these issues, we designed a novel text-based task generation paradigm that elicits unconstrained, intrinsically motivated responses from both humans and LLMs (Fig. 1). Our approach proceeds in two steps.
@@
Second, we compare this human baseline to the output of a widely used LLM (GPT-4o), conditioned on the same value. This allows us to evaluate whether such models can replicate the psychological signatures of human goal-setting or whether fundamental differences remain, particularly in terms of motivational grounding and embodied realism.
@@
Recent advances have positioned LLMs as powerful tools for simulating human cognition and behavior. On one hand, their capabilities in task automation and planning are rapidly improving, pushing the boundaries of what models can execute in complex environments (Park et al. 2023; Yang et al. 2025; Shen et al. 2024; Hu et al. 2024). On the other hand, studies show LLMs can even replicate nuanced human social behaviors, including prosocial irrationality and cognitive biases, suggesting a surprising consistency with human decision-making patterns (Liu et al. 2025). These capabilities have led to growing interest in using LLMs as cognitive models to explore how humans think, decide, and act.
Despite these capabilities, researchers have identified significant gaps in the psychological plausibility of these models. While LLMs can mimic surface-level reasoning and planning, they often lack deeper cognitive mechanisms such as emotion, causality, physical dynamics and social cognition (Hu et al. 2025; Binz and Schulz 2023; Ullman 2024). More fundamentally, because they are trained purely on text, they lack the sensorimotor grounding that shapes human goal formation. Embodied cognition theories emphasize that human goals and reasoning are deeply shaped by our physical experiences and social embeddedness (Varela et al. 2017). This suggests that generating goals rooted in physical interaction and embodied experience is a critical test case for the fidelity of LLM simulators—a test we conduct in this study.
@@ Methods
We conducted two experiments to identify the unique signatures of human task generation by contrasting it with the output of a LLM. Experiment 1 elicited tasks from both human participants and an LLM. For the human baseline, we measured core psychological constructs hypothesized to drive autonomous behavior: stable motivational drivers (personal values, via PVQ21, Schwartz 2021), cognitive strategies for navigating complexity (thinking style, via TWS, Sagiv et al. 2010), along with emotional states as controls. Experiment 2 then involved an independent sample of human raters who evaluated key attributes of the generated tasks, enabling a direct comparison between human and AI outputs.
@@
The experiment employed a 2 (environmental context: high vs. low complexity) $\times$ 2 (social context: social vs. non-social) between-subjects factorial design (Fig. 2B). The environmental context was manipulated to vary the creative constraints. In the high-complexity condition, participants were presented with a list of 32 common living room items. To increase the cognitive demand for task generation in the low-complexity condition, we removed seven items with high entertainment affordances (e.g., Dice, Basketball), leaving 25 common items (e.g., Pens, Chairs).
@@
All object stimuli were selected from the room asset library of the virtual simulation based on Unreal Engine 5 platform for ecological validity.
@@
After completing the psychological scales, participants were randomly assigned to one of the four conditions. They were shown a list of objects and instructed: ”Please imagine that you are in a real room with several items in it. You need to use these items to pass the time.” For each task, participants provided a title, a list of used objects, a description, task setup, goal, and scoring system (Fig. 2A). They reported the mental effort and perceived difficulty for each task generated. Participants were required to generate at least three tasks before choosing to conclude the experiment.
LLM Implementation
We implemented two GPT-4o conditions to test whether the model’s deviation from human goal-setting stems from a simple information deficit or a fundamental mechanism differences. In the raw condition, the model received only the task prompt, serving as a baseline for its default behavior. In the matched condition, each model was given the full psychological profile of a corresponding human participant, including values and cognitive style. This condition directly tested whether the model could translate these psychological constructs into value-driven behavior when provided as semantic data, or whether its lack of internal motivation and embodied grounding imposes a deeper constraint. Both LLMs operated at a temperature of 1.0 to encourage diverse outputs.
@@
The stimulus pool comprised two task sets. To capture the full spectrum of human creativity, we included all 575 valid tasks generated by human participants. For a focused comparison, we compared this to the output from the matched GPT condition, including the first task generated for each of the 176 unique participant profiles. This ensures a representative, non-redundant sample of the model’s performance when conditioned on individual data. Tasks from the raw GPT model were reserved for computational analyses (e.g., topic modeling) but excluded from human rating to streamline the evaluation process. Each of the 75 evaluators then rated a randomized subset of 50 tasks, yielding approximately five independent ratings per task. The evaluation was conducted in a blind manner.
@@
Evaluators rated each task on a 0–10 scale across dimensions capturing creativity (Fun, Novelty) and embodiment (Overall Difficulty, Mental Demand, Physical Demand, Body Part Engagement). These metrics provided a quantitative basis for comparing human- and AI-generated goals (Fig. 2C).
@@
Task Content Analysis.
To identify emergent themes in the generated tasks, we performed topic modeling using BERTopic with embedding model “all-mpnet-base-v2” (Grootendorst 2022; Reimers and Gurevych 2019) on the full corpus of 1718 tasks (575 human, 575 matched GPT, and 568 raw GPT tasks). To measure the creativity of each participant’s output, we computed their task diversity score, defined as the mean cosine distance between the sentence-transformer embeddings of the tasks they generated.
@@
To test our hypotheses about the link between psychological traits and task attributes, we specified a series of linear mixed-effects models (LMMs, Brauer and Curtin 2018) to predict key dependent variables: the rated attributes of Fun, Novelty, and their task content diversity. We first specified a baseline model including only demographic variables and random participants intercepts. Then we specified a main-effects model by adding experimental conditions and psychological predictors like personal values and thinking style. Finally, we included the pre-specified TWS × Environmental Context interaction into the model. The models were compared using a likelihood ratio test (LRT, Moreira 2003) to determine if including the psychological predictors significantly improved model fit.
@@
Figure 3: Personal values and cognitive style shape human goal generation. (A) Regression coefficients of key predictors on task attributes (Novelty, Fun, and Task Diversity). * indicates $p<0.05$, ** indicates $p<0.01$. (B) Openness to Change values predict task fun: as Openness to Change increases, participants’ tasks are rated as more enjoyable. (C) Cognitive style (TWS) interacts with environmental complexity to predict task diversity: in high-complexity environments (black), individuals with intuitive styles (lower TWS) produce more diverse tasks. Bars indicate 95% confidence intervals.
@@ Results
Our analyses proceed in two stages. First, we establish a behavioral signature of human goal generation, demonstrating its systematic link to personal values and environmental context. Second, we contrast this human baseline against the output of LLMs to reveal their gaps.
@@
To establish a human baseline, we first tested how psychological traits predicted people’s goals. Linear mixed-effects models confirmed that task generation was systematically linked to stable personal values (Fig. 3A). The value of Openness to Change significantly predicted higher ratings of both task novelty ($b=0.152$, $t=2.406$, $p=0.017$) and fun (Fig. 3B, $b=0.146$, $t=2.316$, $p=0.022$). In contrast, the self-interest value dimension showed no such relationship. This finding supports the core hypothesis that stable motivational principles guide autonomous goal-setting.
Behavior was not only driven by stable values but was also adapted to the environment. We found that the creativity of generated tasks depended on a significant interaction between an individual’s cognitive style (TWS) and the complexity of the environment (Fig. 3A). This pattern appeared in both subjective ratings of task fun (interaction of TWS and Environmental Context: b=0.140b=-0.140, t=2.640t=-2.640, $p=0.009$), and in objective measures of task diversity (Fig. 3C, b=0.215b=-0.215, t=2.809t=-2.809, $p=0.006$). Specifically, individuals with a more systematic thinking style generated more fun tasks in the simple environment, whereas those with an intuitive style thrived in the complex one. This suggests people flexibly deploy different cognitive strategies to match environmental demands.
Figure 4: Human and GPT generated systematically different goals. (A) Thematic distribution of human and LLM tasks. 1718 tasks generated by human and GPT were clustered into 12 topics and 3 themes: “Physical & Sports Activities”, “Relaxation & Household Activities”, and “Mental & Artistic Activities”. (B) LLMs exhibit a lower propensity for social task generation. (C) LLM tasks are perceived as more mentally and less physically demanding. Black dots stands for the overall difficulty ratings of tasks. Orange and green dots stands for mental and physical load respectively. (D) Engagement ratings of different body parts.
@@
Having established a human behavioral signature, we next compared it to the output of LLMs to identify structural differences. The comparison revealed systematic differences in task content and structure.
@@
We directly compared the distribution of task topics between human participants and LLMs. Topic modeling identified three dominant themes across the dataset: Physical & Sports, Relaxation & Household, and Mental & Artistic. The distribution of tasks across these themes differed significantly between humans and the LLMs ($\chi^{2}(4)=581.45$, $p<0.001$). Human participants generated a balanced mix of activities. In contrast, LLMs showed a bias toward abstract tasks like music and writing (Fig. 4A).
@@
Beyond thematic content, LLM generated significantly fewer social tasks than humans in settings where another person was present ($\chi^{2}(2)=93.19$, $p<0.001$). In these scenarios, 58% of human-generated tasks were multiplayer, with 83% of participants creating at least one social goal. Humans proposed a range of interactive tasks, from simple conversation games like “Guess Who I Am” to competitive games like “Pillow Fight”. In contrast, multiplayer tasks from the Raw and Matched GPT were far less common, accounting for only 29% and 24% of their outputs respectively (Fig. 4B).
@@ A Disembodiment Gap in Perceived Demands
The thematic differences between human and LLM tasks were reflected in their perceived mental and physical demands (Fig. 4C). Tasks generated by the Matched GPT were rated as significantly more mentally demanding (z=11.64z=-11.64, FDR-corrected $p<0.001$) and less physically demanding (z=8.40z=-8.40, $p_{\text{FDR}}<0.001$) than those generated by humans. This aligns with the LLM’s focus on abstract, “Mental & Artistic” activities.
This mental-physical divide was also evident in the specific body parts required for task execution (Fig. 4D). GPT tasks demanded greater brain engagement (z=11.40z=-11.40, $p_{\text{FDR}}<0.001$) but significantly less involvement of the arms (z=5.43z=-5.43, $p_{\text{FDR}}<0.001$) and lower body (z=7.93z=-7.93, $p_{\text{FDR}}<0.001$). Engagement of the eyes, ears, and hands did not differ between the two groups.
@@
Despite their disconnection from physical and social experience, Matched GPT tasks were rated as more novel (z=5.63z=-5.63, $p_{\text{FDR}}<0.001$) and more fun (z=2.73z=-2.73, $p_{\text{FDR}}=0.010$) than those created by humans (Fig. 5). This highlights a key tradeoff: LLMs can generate text that seems creative and engaging, but they do so without grounding in embodied or socially situated experience.